# Force-free wave interaction in magnetar magnetospheres: Computational modeling in axisymmetry

### v1

@@ 2.2 Mode conversion at curved field lines
In this subsection, we consider electric and magnetic fields $\mathbf{E}=\delta\mathbf{E}$ and $\mathbf{B}=\mathbf{B}_{0}+\delta\mathbf{B}$. Here, $\delta\mathbf{E}$ and $\delta\mathbf{B}$ are perturbations to a background field $\mathbf{B}_{0}$ that is not necessarily dipolar. The force-free current on a stationary (i.e., ×$\nabla\times\mathbf{B}_{0}=0$) background expressed in these variables reads (up to second order perturbations):
$\begin{split}&\mathbf{J}_{\rm FFE}\\ &=\left(\nabla\cdot\mathbf{E}\right)\frac{\mathbf{E}\times\mathbf{B}}{B^{2}}+\frac{\mathbf{B}\cdot\left(\nabla\times\mathbf{B}\right)-\mathbf{E}\cdot\left(\nabla\times\mathbf{E}\right)}{B^{2}}\mathbf{B}\\ &=\left(\nabla\cdot\delta\mathbf{E}\right)\frac{\delta\mathbf{E}\times\mathbf{B}_{0}}{B_{0}^{2}}+\frac{\delta\mathbf{B}\cdot\left(\nabla\times\delta\mathbf{B}\right)}{B_{0}^{2}}\mathbf{B}_{0}\\ &\qquad+\frac{\mathbf{B}_{0}\cdot\left(\nabla\times\delta\mathbf{B}\right)}{B_{0}^{2}}\delta\mathbf{B}-\frac{\delta\mathbf{E}\cdot\left(\nabla\times\delta\mathbf{E}\right)}{B_{0}^{2}}\mathbf{B}_{0}\\ &\qquad-\frac{2\mathbf{B}_{0}\cdot\left(\nabla\times\delta\mathbf{B}\right)}{B_{0}^{4}}\left(\delta\mathbf{B}\cdot\mathbf{B}_{0}\right)\mathbf{B}_{0}.\end{split}$ (11)
An Alfvén wave with $\omega=\pm ck_{\parallel}$ in an axisymmetric dipole magnetosphere with $\mathbf{B}_{0}=\mathbf{B}_{\rm dip}$ as described by Equation (6) has a wave vector in the $\left(r,\theta\right)$–plane (a comprehensive analysis of such waves can be found in Chen et al., 2024). The electric perturbation is also in the $\left(r,\theta\right)$–plane and perpendicular to $\mathbf{B}_{\rm dip}$. The magnetic perturbation is along the $\phi$ direction. Their amplitudes are related by $\delta B=-\delta E$. The current defined in Equation (11) has a contribution along the $\phi$ direction with
$\frac{J_{\phi,\rm FFE}}{B_{\rm dip}}=ik_{\perp}\left(\frac{\delta B}{B_{\rm dip}}\right)^{2}.$ (12)
As the Alfvén wave cannot compensate for this current by matching a suitable ×$\nabla\times\delta\mathbf{B}$, a toroidal electric perturbation can be induced. Such perturbations can seed fast magnetosonic waves propagating outwards. The strength of such a fast wave (F) injection depends on the seed Alfvén wave (A) amplitude (with respect to the background field) and its variation $k_{\perp}$ perpendicular to the magnetic field. We can further approximate
$\frac{\delta E_{F}}{B_{\rm dip}}\approx\frac{4\pi}{c}\frac{k_{\perp}}{k_{\parallel}}\left(\frac{\delta B_{A}}{B_{\rm dip}}\right)^{2}\equiv\frac{4\pi}{c}\xi\left(\frac{\delta B_{A}}{B_{\rm dip}}\right),$ (13)
where $\xi=(k_{\perp}/k_{\parallel})(\delta B_{A}/B_{\rm dip})$ is a measure of Alfvén wave (non)-linearity. We analyze the dipolar magnetic fields along the two coordinates
@@
where $\left(r,\theta\right)$ are radius and colatitude of the corresponding spherical coordinate system (Swisdak, 2006; Chen et al., 2022a). The $p$-coordinate labels each field line by its radius on the magnetic equator ($\chi=0$); $q$ varies along the length $s\left(p,\chi\right)$ of the field line. Alfvén waves traveling along the dipole field develop significant shear with a rapid increase of $\xi$. The shear evolution is driven by the gradient
@@
The coupling is strongest at colatitudes in the hemisphere opposite to the injection site. Specifically, Figure 1 has a maximum beyond the equator ($\theta\approx 110^{\circ}$). The coupling strength scales with the wave-driven magnetic stress $b_{\rm eq}^{2}$.
2.3 Non-linear Alfvén wave interaction
@@
Subscripts $1$ and $2$ denote the incoming AWs; unsubscripted quantities refer to the outgoing FMS wave. The coefficient that determines the growth rate of secondary modes of the three-wave interaction is given, for example, in Li et al. (2019, Appendix A)
@@
Here $\delta B_{1}$ and $\delta B_{2}$ are amplitudes for the interacting AWs, and $B_{0}$ is the background magnetic field strength. For two identical waves with reversed polarization $\boldsymbol{k}_{1\perp}=-\boldsymbol{k}_{2\perp}$, this coefficient will vanish; there will be no fast wave produced from the three-wave interaction. With similar arguments than in our derivation of Equation (16), one can infer a significant spatial dependence of the interaction coefficient: off-equatorial interactions or asymmetric wave injection scenarios could change – likely decrease – FMS wave coupling. The solution space to the kinetic condition (Equation 17) depends on the resolved dimensions and the allowed wave vectors $\mathbf{k}$. Capturing space in three dimensions will allow more orientations of the propagating waves and, thus, more resonant channels than in two dimensions. Therefore, considering three-dimensional setups of the vast parameter space in future work may lead to different conversion dynamics (including relativistic turbulence; see, e.g., Li et al., 2019; Ripperda et al., 2021; TenBarge et al., 2021). We discuss the implications and limitations of dimensionality in more detail in Section 4.2.
@@ 3.1 Setup
All our simulations initialize a dipolar magnetic field anchored to a perfect conductor of radius $R_{\ast}$ following Equation (6). We use spherical coordinates with axisymmetry (similar to Mahlmann & Aloy, 2021, on the same infrastructure). The computational domain encloses the area $(r,\theta)\in\left[R_{\ast},40R_{\ast}\right]\times\left[0,\pi\right]$ with uniform spacing. We vary the resolution by changing the number of grid points in each direction, with $N_{r}\in\left\{4992,9984,19968,39936\right\}$ and $N_{\theta}=\left[400,800,1600,3200\right]$. We also measure the radial resolution by the number of grid points per stellar radius $N_{R_{\ast}}\in\left\{128,256,512,1024\right\}$. Beyond $r/R_{\ast}=40$, the radial grid spacing increases by a factor $a=1.001$ in each cell, effectively allowing a logarithmic spacing up to very large radii. We integrate this system with seventh-order accurate spatial reconstruction (MP7, Suresh & Huynh, 1997) and $\text{CFL}\approx 0.2$.
@@ 3.2 Wave trains ‘bounce’ in the inner magnetosphere
Figure 3: Outgoing FMS wave luminosity injected by long-envelope AWs (see Figure LABEL:fig:a1bounce) measured at $r/R_{\ast}=5$ and normalized to the injected AW energy $E_{\rm inj}$. We vary the injection amplitude for a fixed numerical resolution. The inset plot shows the peak luminosity as a function of the wave amplitude $b_{\rm eq}$, following a trend $\propto b_{\rm eq}^{2}$.
@@
We first fix the injection amplitude such that $b_{\rm eq}\approx 0.2$. Figure LABEL:fig:a1bounce (and animation in the supplementary material; Mahlmann et al., 2024a) shows an Alfvén wave envelope propagating through the inner magnetosphere and building up significant shear (panel A). The sheared AWs inject FMS waves (panel a) as described in Section 3.3. Once the AWs reach the stellar surface of the opposite hemisphere, they bounce off it and propagate back (panel B/C). FMS waves with amplitudes of opposite signs are produced upon each AW crossing (panels b/c/d). The bottom panels of Figure LABEL:fig:a1bounce display short-wavelength modulations of the generated FMS waves. Fourier analysis of the modulation shows that their wavelength is half of the wavelength of the injected Alfvén waves (see Appendix B). The high-frequency variations of seed AWs only imprint on the outgoing FMS waves with twice the AW frequency during the first magnetospheric crossing. Later FMS waves have a much longer wavelength comparable to the timescale $T_{\rm bounce}$ of AWs propagating through the magnetosphere. For the chosen injection colatitude of $\theta_{s}=45^{\circ}$, field lines have a length of $L\approx 3.5R_{\ast}$. The bottom panels of Figure LABEL:fig:a1bounce illustrate the growth of FMS wave amplitude relative to the background magnetic field. Relatively small FMS production amplitudes ($\delta E_{F}/B\lesssim 0.05$) can grow to significant relative field strengths (with total field strengths $E\approx B$) within a few tens of stellar radii (see Figure LABEL:fig:a1bounce, panel d, and Equation 10). We note that, in this example, the electric zones develop only for the second half of the first generated FMS wave packet, the preceding high-frequency component of opposite $\delta E_{F}$ polarity propagates unobstructed. In Figure 3, we analyze how AWs with increasing amplitudes $b_{\rm eq}$ translate into a wave Poynting flux $\mathbf{S}=\delta\mathbf{E}\times\delta\mathbf{B}$ associated to outgoing FMS waves. Integrated over a spherical surface, we obtain the luminosity $L_{r}$. The peak luminosity measured at $r/R_{\ast}=5$ scales with the square of the AW amplitude. Equation (13) predicts this enhancement with the magnetic stress carried by the AW. The peaks of the luminosity are separated by $ct/R_{\ast}\approx 2$ (see further discussion below).
@@
Figure 4: Resolution requirements for late-stage FMS wave generation. The bottom panel is as Figure 3 but for a fixed injection amplitude ($b_{\rm eq}\approx 0.2$) and varying numerical resolution. The top panel shows the excess magnetospheric energy $\Delta E_{\rm tot}=E_{\rm tot}-E_{\rm dip}$, normalized to the injected AW energy $E_{\rm inj}$. For an injection colatitude of $\theta_{s}=45^{\circ}$, field lines have a length of $L\approx 3.5R_{\ast}$; we denote the light-crossing time of this length as $T_{\rm bounce}$.
@@
The strong shear that AWs experience (especially after several magnetospheric crossings) makes them susceptible to grid diffusion. Therefore, it is instructive to probe the numerical convergence of the FMS wave injection. Figure 4 (bottom panel) shows the luminosity of outgoing FMS waves measured at $r/R_{\ast}=5$ for varying resolutions $N_{R_{\ast}}\in\left\{128,256,512,1024\right\}$. Higher resolutions capture the sheared waves significantly better. The top panel of Figure 4 tracks the evolution of the excess energy $\Delta E_{\rm tot}=E_{\rm tot}-E_{\rm dip}$, which corresponds to the energy injected in AWs initially, and to the combined energy in AWs and FMS waves at later times $t\gtrsim T$. For $N_{R_{\ast}}=128$, the entire injected AW energy dissipates right after the first magnetospheric ‘bounce’ of duration $T_{\rm bounce}$ (here, this scale denotes the time it takes a seed AW to cross the magnetosphere along a dipolar field line). Higher resolutions delay the decay of excess energy to later times. Hence, the decay is driven by numerical diffusion as the energy cascades down to scales of the grid resolution. The highest resolution of $N_{R_{\ast}}=1024$ allows for several magnetospheric bounces with mild diffusion at the grid scale. Our results show that all the chosen resolutions capture the initial phase of FMS wave production equally well. Only after the second magnetospheric crossing and $T_{\rm bounce}\gtrsim 2$, the FMS wave energy significantly deviates for the lower resolution setups. The total efficiency of FMS wave conversion ranges between $0.23\%$ ($N_{R_{\ast}}=128$), $0.23\%$ ($N_{R_{\ast}}=256$), $0.28\%$ ($N_{R_{\ast}}=512$), and $0.30\%$ ($N_{R_{\ast}}=1024$). For these parameters, the FMS energy varies by $30\%$, attributed purely to the numerical resolution capturing (or not) the strong shear of Alfvén waves in the decaying tail of FMS wave pulses. In Appendix A, we estimate that $0.16\%$ of the injected AW energy (around half of the entire energy transferred to FMS waves) is converted to FMS waves in the tail after $ct/R_{\ast}\approx 8.43$. Focusing only on this period emphasizes the damping effect of grid diffusion for late times [^3] We use this comparison to an approximate ‘true’ amount of energy stored in the tail (Appendix A) to estimate the order of convergence achieved with our method in resolving the decaying FMS wave tail. For $N_{R_{\ast}}\gtrsim 256$, we empirically find a convergence order of $1.27$. This is significantly below the theoretical convergence order of our method (seventh order for smooth regions, cf. Mahlmann et al., 2020b). The break-down of convergence order is a strong indication of the difficulties in resolving the ongoing force-free wave interactions in the inner magnetar magnetosphere.: the highest probed resolution ($N_{R_{\ast}}=1024$) recovers around $90\%$ of the tail energy $\tilde{E}_{\rm tail}$, the lowest probed resolution ($N_{R_{\ast}}=128$) only recovers $38\%$ and fails to resolve the tail modulation with $T_{\rm bounce}$.

### v2

@@ 2.2 Mode conversion at curved field lines
In this subsection, we consider electric and magnetic fields $\mathbf{E}=\delta\mathbf{E}$ and $\mathbf{B}=\mathbf{B}_{0}+\delta\mathbf{B}$. Here, $\delta\mathbf{E}$ and $\delta\mathbf{B}$ are perturbations to a background field $\mathbf{B}_{0}$ that is not necessarily dipolar. The force-free current on a stationary (i.e., ×$\nabla\times\mathbf{B}_{0}=0$) background expressed in these variables reads (up to second order-perturbations) as
$\begin{split}&\mathbf{J}_{\rm FFE}\\ &=\left(\nabla\cdot\mathbf{E}\right)\frac{\mathbf{E}\times\mathbf{B}}{B^{2}}+\frac{\mathbf{B}\cdot\left(\nabla\times\mathbf{B}\right)-\mathbf{E}\cdot\left(\nabla\times\mathbf{E}\right)}{B^{2}}\mathbf{B}\\ &=\left(\nabla\cdot\delta\mathbf{E}\right)\frac{\delta\mathbf{E}\times\mathbf{B}_{0}}{B_{0}^{2}}+\frac{\delta\mathbf{B}\cdot\left(\nabla\times\delta\mathbf{B}\right)}{B_{0}^{2}}\mathbf{B}_{0}\\ &\qquad+\frac{\mathbf{B}_{0}\cdot\left(\nabla\times\delta\mathbf{B}\right)}{B_{0}^{2}}\delta\mathbf{B}-\frac{\delta\mathbf{E}\cdot\left(\nabla\times\delta\mathbf{E}\right)}{B_{0}^{2}}\mathbf{B}_{0}\\ &\qquad-\frac{2\mathbf{B}_{0}\cdot\left(\nabla\times\delta\mathbf{B}\right)}{B_{0}^{4}}\left(\delta\mathbf{B}\cdot\mathbf{B}_{0}\right)\mathbf{B}_{0}\\ &\qquad+\frac{\mathbf{B}_{0}\cdot\left(\nabla\times\delta\mathbf{B}\right)}{B_{0}^{2}}\mathbf{B}_{0}.\end{split}$ (11)
An Alfvén wave with $\omega=\pm ck_{\parallel}$ in an axisymmetric dipole magnetosphere with $\mathbf{B}_{0}=\mathbf{B}_{\rm dip}$ as described by Equation (6) has a wavevector in the $\left(r,\theta\right)$–plane (a comprehensive analysis of such waves can be found in Chen et al., 2024). The electric perturbation is also in the $\left(r,\theta\right)$–plane and perpendicular to $\mathbf{B}_{\rm dip}$. The magnetic perturbation is along the $\phi$ direction. Their amplitudes are related by $\delta B=-\delta E$. Analysis of wave-like perturbations shows that if the current defined in Equation (11) has a second-order contribution along the $\phi$ direction (due to curvature effects such that the first three terms do not cancel out as for straight field lines, see Chen et al., 2024), it has the dependence
$\frac{J_{\phi,\rm FFE}}{B_{\rm dip}}\propto ik_{\perp}\left(\frac{\delta B}{B_{\rm dip}}\right)^{2}.$ (12)
If the Alfvén wave cannot compensate for this current by matching a suitable ×$\nabla\times\delta\mathbf{B}$, a toroidal electric perturbation can be induced. Such perturbations can seed FMS propagating outward. The strength of such a fast wave (F) injection can be roughly constrained by analyzing Equation (3) in the limit of no ×$\nabla\times\delta\mathbf{B}$ compensation. It then depends on the seed Alfvén wave (A) amplitude (with respect to the background field) and its variation $k_{\perp}$ perpendicular to the magnetic field. We can further approximate
$\frac{\delta E_{F}}{B_{\rm dip}}\propto\frac{4\pi}{c}\frac{k_{\perp}}{k_{\parallel}}\left(\frac{\delta B_{A}}{B_{\rm dip}}\right)^{2}\equiv\frac{4\pi}{c}\xi\left(\frac{\delta B_{A}}{B_{\rm dip}}\right),$ (13)
where $\xi=(k_{\perp}/k_{\parallel})(\delta B_{A}/B_{\rm dip})$ is a measure of Alfvén wave (non)linearity. We analyze the dipolar magnetic fields along the two coordinates
@@
where $\left(r,\theta\right)$ are the radius and colatitude of the corresponding spherical coordinate system (Swisdak, 2006; Chen et al., 2022a). The $p$-coordinate labels each field line by its radius on the magnetic equator ($\chi=0$); $q$ varies along the length $s\left(p,\chi\right)$ of the field line. Alfvén waves traveling along the dipole field develop significant shear with a rapid increase of $\xi$. The shear evolution is driven by the gradient
@@
The coupling is strongest at colatitudes in the hemisphere opposite to the injection site. Specifically, Figure 1 has a maximum beyond the equator ($\theta\approx 110^{\circ}$). The coupling strength scales with the wave-driven magnetic stress $b_{\rm eq}^{2}$. We note that the fast wave properties outlined in this section are only rough estimates of the expected trends. We ultimately rely on fully fledged FFE simulations (Section 3) to solve the nonlinear field evolution and consider the complex field line geometry.
2.3 Nonlinear Alfvén wave interaction
@@
Subscripts $1$ and $2$ denote the incoming Alfvén waves; unsubscripted quantities refer to the outgoing FMS wave. The coefficient that determines the growth rate of secondary modes of the three-wave interaction is given, for example, in Li et al. (2019, Appendix A)
@@
Here, $\delta B_{1}$ and $\delta B_{2}$ are amplitudes for the interacting Alfvén waves, and $B_{0}$ is the background magnetic field strength. For two identical waves with reversed polarization $\boldsymbol{k}_{1\perp}=-\boldsymbol{k}_{2\perp}$, this coefficient will vanish; there will be no fast wave produced from the three-wave interaction. With similar arguments than in our derivation of Equation (16), one can infer a significant spatial dependence of the interaction coefficient: off-equatorial interactions or asymmetric wave injection scenarios could change – likely decrease – FMS wave coupling. The solution space to the kinetic condition (Equation 17) depends on the resolved dimensions and the allowed wavevectors $\mathbf{k}$. Capturing space in three dimensions will allow more orientations of the propagating waves and, thus, more resonant channels than in two dimensions. Therefore, considering three-dimensional setups of the vast parameter space in future work may lead to different conversion dynamics (including relativistic turbulence; see, e.g., Li et al., 2019; Ripperda et al., 2021; TenBarge et al., 2021). We discuss the implications and limitations of dimensionality in more detail in Section 4.2.
@@ 3.1 Setup
All our simulations initialize a dipolar magnetic field anchored to a perfect conductor of radius $R_{\ast}$ following Equation (6). We use spherical coordinates with axisymmetry (similar to Mahlmann & Aloy, 2021, on the same infrastructure). The computational domain encloses the area $(r,\theta)\in\left[R_{\ast},40R_{\ast}\right]\times\left[0,\pi\right]$ with uniform spacing. We vary the resolution by changing the number of grid points in each direction, with $N_{r}\in\left\{4992,9984,19968,39936\right\}$ and $N_{\theta}=\left[400,800,1600,3200\right]$. We also measure the radial resolution by the number of grid points per stellar radius $N_{R_{\ast}}\in\left\{128,256,512,1024\right\}$. Beyond $r/R_{\ast}=40$, the radial grid spacing increases by a factor $a=1.001$ in each cell, effectively allowing a logarithmic spacing up to very large radii. We integrate this system with seventh-order accurate spatial reconstruction (MP7; Suresh & Huynh, 1997) and $\text{CFL}\approx 0.2$.
@@ 3.2 Wave trains ‘bounce’ in the inner magnetosphere
Figure 3: Outgoing FMS wave luminosity injected by long-envelope Alfvén waves (see Figure LABEL:fig:a1bounce) measured at $r/R_{\ast}=5$ and normalized to the injected Alfvén wave energy $E_{\rm inj}$. We vary the injection amplitude for a fixed numerical resolution. The inset plot shows the peak luminosity as a function of the wave amplitude $b_{\rm eq}$, following a trend $\propto b_{\rm eq}^{2}$.
@@
We first fix the injection amplitude such that $b_{\rm eq}\approx 0.2$. Figure LABEL:fig:a1bounce (and animation in the supplementary material; Mahlmann et al., 2024a) shows an Alfvén wave envelope propagating through the inner magnetosphere and building up significant shear (panel A). The sheared Alfvén waves inject FMS waves (panel a) as described in Section 3.3. Once the Alfvén waves reach the stellar surface of the opposite hemisphere, they bounce off it and propagate back (panel B/C). FMS waves with amplitudes of opposite signs are produced upon each Alfvén wave crossing (panels b/c/d). The bottom panels of Figure LABEL:fig:a1bounce display short-wavelength modulations of the generated FMS waves. Fourier analysis of the modulation shows that their wavelength is half of the wavelength of the injected Alfvén waves (see Appendix B). The high-frequency variations of seed Alfvén waves only imprint on the outgoing FMS waves with twice the Alfvén wave frequency during the first magnetospheric crossing. Later FMS waves have a much longer wavelength comparable to the timescale $T_{\rm bounce}$ of Alfvén waves propagating through the magnetosphere. For the chosen injection colatitude of $\theta_{s}=45^{\circ}$, field lines have a length of $L\approx 3.5R_{\ast}$. The bottom panels of Figure LABEL:fig:a1bounce illustrate the growth of FMS wave amplitude relative to the background magnetic field. Relatively small FMS production amplitudes ($\delta E_{F}/B\lesssim 0.05$) can grow to significant relative field strengths (with total field strengths $E\approx B$) within a few tens of stellar radii (see Figure LABEL:fig:a1bounce, panel d, and Equation 10). We note that, in this example, the electric zones develop only for the second half of the first generated FMS wave packet, the preceding high-frequency component of opposite $\delta E_{F}$ polarity propagates unobstructed. In Figure 3, we analyze how Alfvén waves with increasing amplitudes $b_{\rm eq}$ translate into a wave Poynting flux $\mathbf{S}=\delta\mathbf{E}\times\delta\mathbf{B}$ associated to outgoing FMS waves. Integrated over a spherical surface, we obtain the luminosity $L_{r}$. The peak luminosity measured at $r/R_{\ast}=5$ scales with the square of the Alfvén wave amplitude. Equation (13) predicts this enhancement with the magnetic stress carried by the AW. The peaks of the luminosity are separated by $ct/R_{\ast}\approx 2$ (see further discussion below).
@@
Figure 4: Resolution requirements for late-stage FMS wave generation. The bottom panel is as Figure 3 but for a fixed injection amplitude ($b_{\rm eq}\approx 0.2$) and varying numerical resolution. The top panel shows the excess magnetospheric energy $\Delta E_{\rm tot}=E_{\rm tot}-E_{\rm dip}$, normalized to the injected Alfvén wave energy $E_{\rm inj}$. For an injection colatitude of $\theta_{s}=45^{\circ}$, field lines have a length of $L\approx 3.5R_{\ast}$; we denote the light-crossing time of this length as $T_{\rm bounce}$.
@@
The strong shear that Alfvén waves experience (especially after several magnetospheric crossings) makes them susceptible to grid diffusion. Therefore, it is instructive to probe the numerical convergence of the FMS wave injection. Figure 4 (bottom panel) shows the luminosity of outgoing FMS waves measured at $r/R_{\ast}=5$ for varying resolutions $N_{R_{\ast}}\in\left\{128,256,512,1024\right\}$. Higher resolutions capture the sheared waves significantly better. The top panel of Figure 4 tracks the evolution of the excess energy $\Delta E_{\rm tot}=E_{\rm tot}-E_{\rm dip}$, which corresponds to the energy injected in Alfvén waves initially, and to the combined energy in Alfvén waves and FMS waves at later times $t\gtrsim T$. For $N_{R_{\ast}}=128$, the entire injected Alfvén wave energy dissipates right after the first magnetospheric ‘bounce’ of duration $T_{\rm bounce}$ (here, this scale denotes the time it takes a seed Alfvén wave to cross the magnetosphere along a dipolar field line). Higher resolutions delay the decay of excess energy to later times. Hence, the decay is driven by numerical diffusion as the energy cascades down to scales of the grid resolution. The highest resolution of $N_{R_{\ast}}=1024$ allows for several magnetospheric bounces with mild diffusion at the grid scale. Our results show that all the chosen resolutions capture the initial phase of FMS wave production equally well. Only after the second magnetospheric crossing and $T_{\rm bounce}\gtrsim 2$, the FMS wave energy significantly deviates for the lower resolution setups. The total efficiency of FMS wave conversion ranges between $0.23\%$ ($N_{R_{\ast}}=128$), $0.23\%$ ($N_{R_{\ast}}=256$), $0.28\%$ ($N_{R_{\ast}}=512$), and $0.30\%$ ($N_{R_{\ast}}=1024$). For these parameters, the FMS energy varies by $30\%$, attributed purely to the numerical resolution capturing (or not) the strong shear of Alfvén waves in the decaying tail of FMS wave pulses. In Appendix A, we estimate that $0.16\%$ of the injected Alfvén wave energy (around half of the entire energy transferred to FMS waves) is converted to FMS waves in the tail after $ct/R_{\ast}\approx 8.43$. Focusing only on this period emphasizes the damping effect of grid diffusion for late times [^3] We use this comparison to an approximate ‘true’ amount of energy stored in the tail (Appendix A) to estimate the order of convergence achieved with our method in resolving the decaying FMS wave tail. For $N_{R_{\ast}}\gtrsim 256$, we empirically find a convergence order of $1.27$. This is significantly below the theoretical convergence order of our method (seventh order for smooth regions, cf. Mahlmann et al., 2020b). The break-down of convergence order is a strong indication of the difficulties in resolving the ongoing force-free wave interactions in the inner magnetar magnetosphere.: the highest probed resolution ($N_{R_{\ast}}=1024$) recovers around $90\%$ of the tail energy $\tilde{E}_{\rm tail}$, the lowest probed resolution ($N_{R_{\ast}}=128$) only recovers $38\%$ and fails to resolve the tail modulation with $T_{\rm bounce}$.